# Performance bounds for priority-based stochastic coflow scheduling

### v1

@@ 6 Concluding remarks
We addressed the coflow scheduling problem in the stochastic setting, assuming that the objective is to minimize the weighted expected completion time of coflows. We proposed a simple priority policy in which the priority order is computed with $\mathtt{Sincronia}$using the expected processing times instead of the unknown actual processing times. Performance guarantees are established for this policy using a polyhedral relaxation of the performance space of stochastic coflow scheduling policies, first for arbitrary probability distributions of processing times, and then for some specific distributions. Empirical results show that this priority policy performs very well in practice, and much better than the theoretical upper bounds suggest.
We see the following possible future directions emanating from this work. First, we plan to design preemptive scheduling policies leveraging the information on the attained service of coflows. Another interesting direction will be to propose policies that choose between RR and NC based on the variability of the flow-sizes. Further, it will also be interesting to compute the approximation ratio for a richer class of policies than the order-based ones in this paper. As we saw, certain policies like RR as not included in this class. Finally, we intend to establish tighter bounds for some other specific distributions of processing times, and hope to propose methods for computing the mean CCT in some asymptotic regimes such as large number of coflows and ports.
@@ Appendix A Proof of Theorem 1
Let $r$ be a work-conserving and nonanticipative rate-allocation policy and define the random variable $Z_{\ell,k}=C_{k}(r)-P_{\ell,k}$. Consider any fixed realization of flow volumes. The latter yields a specific realization of the processing times $p_{\ell,k}$ and of the completion times $c_{k}(r)$ (we use lowercase letters to denote the particular values that our random variables take for this specific realization of flow volumes). Assume that, for this realization, coflows complete in the order $\pi$, that is, $c_{\pi(1)}(r)\leq c_{\pi(2)}(r)\leq\ldots\leq c_{\pi(n)}(r)$. As coflow $\pi(k)$ completes after all coflows $\pi(j)$, $j=1,\ldots,k-1$, we clearly have $c_{\pi(k)}(r)\geq\sum_{j\leq k}p_{\ell,\pi(j)}$ for all ports $\ell$. It yields $z_{\ell,\pi(k)}\geq\sum_{j<k}p_{\ell,\pi(j)}\delta_{\pi(j)}^{A}$ for any subset $A\subseteq{\mathcal{C}}$ , where $\delta_{\pi(j)}^{A}$ is $1$ if $\pi(j)\in A$ and 00 otherwise. We thus obtain
@@ Appendix B Proof of Lemma 2
$\leq(1-\gamma\mu_{min}t)^{-\frac{\mu_{\ell,j}}{\gamma\mu_{min}}},\quad t<\gamma\mu_{min}^{-1},$ (26)
@@
Note that the LHS is an increasing function of $z$ and any $z$ satisfying the equality is larger than $1$. Applying the bound [30]
@@ Appendix C Proof of Lemma 3
where inequality (29) follows from Jensen’s inequality and eqality (33) follows from the independence assumption. As we assume a normal distribution for the $P_{\ell,j}$, we have $\mathbb{E}\left[exp(tP_{\ell,j})\right]=exp\left(t\mu_{\ell,j}+\frac{\sigma_{\ell,j}^{2}}{2}t^{2}\right)$, where $\sigma_{\ell,j}^{2}=CV_{\ell,j}^{2}\mu_{\ell,j}^{2}$ is the variance. It yields
@@
$\mathbb{E}\left[\max_{\ell}\sum_{j=1}^{k}P_{\ell,j}\right]\leq\max_{\ell}\sum_{j=1}^{k}\mu_{\ell,j}+\sqrt{2\log(L)}\max_{\ell}\sqrt{\sum_{j=1}^{k}CV_{\ell,j}^{2}\mu_{\ell,j}^{2}}.$ (34)
@@
$\mathbb{E}\left[\max_{\ell}\sum_{j=1}^{k}P_{\ell,j}\right]$ $\leq\max_{\ell}\sum_{j=1}^{k}\mu_{\ell,j}+\sqrt{2\log(L)}\,CV_{max}\,\max_{\ell}\sqrt{\sum_{j=1}^{k}\mu_{\ell,j}^{2}},$ (35)
$\leq\max_{\ell}\sum_{j=1}^{k}\mu_{\ell,j}+\sqrt{2\log(L)}\,CV_{max}\,\max_{\ell}\sum_{j=1}^{k}\mu_{\ell,j},$ (36)
so that $\alpha\leq 1+\sqrt{2\log(L)}\,CV_{max}$, as claimed. ∎

### v2

@@ 6 Concluding remarks
We addressed the coflow scheduling problem in the stochastic setting, assuming that the objective is to minimize the weighted expected completion time of coflows. We proposed a simple priority policy in which the priority order is computed with $\mathtt{Sincronia}\$using the expected processing times instead of the unknown actual processing times. Performance guarantees are established for this policy using a polyhedral relaxation of the performance space of stochastic coflow scheduling policies, first for arbitrary probability distributions of processing times, and then for some specific distributions. Empirical results show that this priority policy performs very well in practice, and much better than the theoretical upper bounds suggest.
We see the following possible future directions emanating from this work. First, we plan to design preemptive scheduling policies leveraging the information on the attained service of coflows. Another interesting direction will be to propose policies that choose between RR and NC based on the variability of the flow-sizes. Further, it will also be interesting to compute the approximation ratio for a richer class of policies than the order-based ones in this paper. As we saw, certain policies like RR are not included in this class. Finally, we intend to establish tighter bounds for some other specific distributions of processing times, and hope to propose methods for computing the mean CCT in some asymptotic regimes such as large number of coflows and ports.
@@ Appendix A Proof of Theorem 1
Let $r$ be a work-conserving and nonanticipative rate-allocation policy and define the random variable $Z_{\ell,k}=C_{k}(r)-P_{\ell,k}$. Consider any fixed realization of flow volumes. The latter yields a specific realization of the processing times $p_{\ell,k}$ and of the completion times $c_{k}(r)$ (we use lowercase letters to denote the particular values that our random variables take for this specific realization of flow volumes). Assume that, for this realization, coflows complete in the order $\pi$, that is, $c_{\pi(1)}(r)\leq c_{\pi(2)}(r)\leq\ldots\leq c_{\pi(n)}(r)$. As coflow $\pi(k)$ completes after all coflows $\pi(j)$, $j=1,\ldots,k-1$, we clearly have $c_{\pi(k)}(r)\geq\sum_{j\leq k}p_{\ell,\pi(j)}$ for all ports $\ell$. It yields $z_{\ell,\pi(k)}\geq\sum_{j<k}p_{\ell,\pi(j)}\delta_{\pi(j)}^{A}$ for any subset $A\subseteq{\mathcal{C}}$ , where $\delta_{\pi(j)}^{A}$ is $1$ if $\pi(j)\in A$ and 0 otherwise. We thus obtain
@@ Appendix B Proof of Lemma 2
$\leq(1-\gamma\mu_{min}t)^{-\frac{\mu_{\ell,j}}{\gamma\mu_{min}}},\quad t<(\gamma\mu_{min}^{-1}),$ (27)
@@
Note that the LHS is an increasing function of $z$ and any $z$ satisfying the equality is larger than $1$. Applying the bound [T07]
@@ Appendix C Proof of Lemma 3
where inequality (30) follows from Jensen’s inequality and eqality (34) follows from the independence assumption. As we assume a normal distribution for the $P_{\ell,j}$, we have $\mathbb{E}\left[exp(tP_{\ell,j})\right]=exp\left(t\mu_{\ell,j}+\frac{\sigma_{\ell,j}^{2}}{2}t^{2}\right)$, where $\sigma_{\ell,j}^{2}=\eta_{\ell,j}^{2}\mu_{\ell,j}^{2}$ is the variance. It yields
@@
$\mathbb{E}\left[\max_{\ell}\sum_{j=1}^{k}P_{\ell,j}\right]\leq\max_{\ell}\sum_{j=1}^{k}\mu_{\ell,j}+\sqrt{2\log(L)}\max_{\ell}\sqrt{\sum_{j=1}^{k}\eta_{\ell,j}^{2}\mu_{\ell,j}^{2}}.$ (35)
@@
$\mathbb{E}\left[\max_{\ell}\sum_{j=1}^{k}P_{\ell,j}\right]$ $\leq\max_{\ell}\sum_{j=1}^{k}\mu_{\ell,j}+\sqrt{2\log(L)}\,\eta_{max}\,\max_{\ell}\sqrt{\sum_{j=1}^{k}\mu_{\ell,j}^{2}},$ (36)
$\leq\max_{\ell}\sum_{j=1}^{k}\mu_{\ell,j}+\sqrt{2\log(L)}\,\eta_{max}\,\max_{\ell}\sum_{j=1}^{k}\mu_{\ell,j},$ (37)
so that $\alpha\leq 1+\sqrt{2\log(L)}\,\eta_{max}$, as claimed. ∎